There are ten pieces of pizza. Anthony ate 9/20 of the pizza for dinner. He ate 1/16 of the pizza for a bedtime snack. How much of the pizza has he eaten in all?
answer as a fraction, reduced to lowest terms.

Answer by rfer(16322)

9/20 + 1/16=
36/80 + 5 80 = 41/80